# Time arrow in open-boundary one-dimensional stochastic dynamics

### v1

@@ I Introduction
The nonequilibrium nature in similar temperature-gradient systems has been discussed in Refs. [8, 17, 18]. These studies revealed the discrepancy in overdamped stochastic dynamics, as well as its shortage of contribution in total entropy dissipation of temperature-gradient systems. For example, in the work of Celani et al.[17], they obtained in underdamped stochastic dynamics an additional source of time irreversibility, or “entropy anomaly” in overall dissipation. This irreversible behavior is attributed to the circulating steady-state behavior in the phase space, and such an entropy anomaly persists even if one proceeds towards the overdamped limit. In our current work, we consider the overdamped stochastic dynamics only, and we find yet another source of time irreversibility, which can arise due to the finiteness of time intervals between successive thermal kicks. And in contrast to the work in [17], which assumes a smooth temperature variation over space, we choose to study the other extreme, in which temperature changes abruptly.
@@ II System
Figure 1: Illustrations of two scenarios. Arrows represent probability currents. Points $C$, $B$, $A$ represent successive discrete positions from left to right, while point $A$ is at a higher temperature. (a) The chain-connecting scenario. Note that the net steady state current between nodes has to be zero everywhere (detailed balance). (b) The loop scenario. In finite-timestep stochastic dynamics, point $A$ can reach $C$ directly over one single timestep $\Delta t$, thanks to the large diffusion parameter in $A$. This fact gives rise to the loop-like scenario and can contribute to a virtually gyrating current at the steady state.
@@ IV.1 Joint Transition Probability
Figure 2: Ratio of joint probabilities between forward and backward transitions. The counting region width $\delta$ is chosen to be identical to the global root-mean-square (RMS) displacement of the particle dynamics. Blue dots: result where joint probabilities were extracted directly from simulation data; green dots: result evaluated from Eq. (14), via the NESS probability distribution $P(x)$ from simulation.
@@ IV.2 NESS probability distribution
Figure 3: Comparison of NESS probability distributions $P(x)$ for a Brownian particle in a harmonic potential with inhomogeneous temperature profile. Blue dots represent numerical simulation results which the simulation parameters are described in section IV, part A. The orange solid line shows the theoretical prediction from Eq. (15) using the step-function temperature profile, which exhibits a sharp discontinuity at $x_{0}$. The green dashed line represents the modified theoretical prediction using the smooth temperature profile from Eq. (17) (see inset), demonstrating a better agreement with simulation results, particularly in the vicinity of the temperature discontinuity.
@@ IV.3 Study in NESS probability current
Figure 4: Probability current $j(x)$, computed via $j=P(x)v(x)$, where $v(x)$ is obtained through finite difference approximation of particle velocities. Note that for one-dimensional open systems, the ideal probability current has to be zero at all positions.
@@ V Modified Temperature Profile and Probability Distribution
In Fig. 3, we find that, with the use of the effective temperature profile, the result of FPE analysis fit remarkably well with the simulation result in probability distribution. We have compared the simulation results with the FPE analysis using the effective temperature profile over various timesteps and temperature settings, while both results stay in pretty good agreement with each other.
The above comparison suggests that one may still represent a one-dimensional stochastic simulation over discrete timesteps via an FPE analysis rather faithfully. On the other hand, as we noted earlier, any one-dimensional Fokker-Planck-type equation must obey the detailed-balance or reversible steady-state dynamics, thanks to the chain-connecting structure of its underlying model. While the latter feature is inconsistent with our simulation observation, we remark here that this paradox can be resolved, knowing that time irreversibility is implicitly incorporated through the effective temperature formalism. To be more specific, an effective temperature at a particular position includes information of the temperature at its last timestep, and such a memory effect is in fact irreversible.

### v2

@@ I Introduction
The nonequilibrium nature in similar temperature-gradient systems has been discussed in Refs. [8, 17, 18, 19]. These studies revealed the discrepancy in overdamped stochastic dynamics, as well as its shortage of contribution in total entropy dissipation of temperature-gradient systems. For example, in the work of Celani et al.[18], they obtained in underdamped stochastic dynamics an additional source of time irreversibility, or “entropy anomaly” in overall dissipation. This irreversible behavior is attributed to the circulating steady-state behavior in the phase space, and such an entropy anomaly persists even if one proceeds towards the overdamped limit. In our current work, we consider the overdamped stochastic dynamics only, and we find yet another source of time irreversibility, which can arise due to the finiteness of time intervals between successive thermal kicks. And in contrast to the work in [18], which assumes a smooth temperature variation over space, we choose to study the other extreme, in which temperature changes abruptly.
@@ II System
Figure 1: Illustrations of two scenarios. Arrows represent probability currents. Points $C$, $B$, $A$ represent successive discrete positions from left to right, while point $A$ is at a higher temperature. (a) The chain-connecting scenario. Note that the net steady state current between nodes has to be zero everywhere (detailed balance). (b) The loop scenario. In finite-timestep stochastic dynamics, point $A$ can reach $C$ directly over one single timestep $\Delta t$, thanks to the large diffusion parameter in $A$. This feature gives rise to the loop-like scenario and can contribute to a virtually gyrating current at the steady state.
@@ IV.1 Joint Transition Probability
Figure 2: Ratio of joint probabilities between forward and backward transitions. The counting region width $\delta$ is determined by the global root-mean-square (RMS) displacement of the Brownian particle. Blue dots: result where joint probabilities were extracted directly from simulation data; green dots: result evaluated from Eq. (14), via the NESS probability distribution $P(x)$ from simulation.
@@ IV.2 NESS probability distribution
Figure 3: Comparison of NESS probability distributions $P(x)$ for a Brownian particle in a harmonic potential with inhomogeneous temperature profile. Blue dots represent the simulation results; the orange solid line shows the theoretical prediction from Eq. (15) using the step-function temperature profile; the green dashed line represents the modified theoretical prediction using the effective temperature profile from Eq. (17) (see inset).
@@ IV.3 Study in NESS probability current
Figure 4: Probability current $j(x)$, evaluated via $j=P(x)v(x)$, where $v(x)$ is obtained through finite difference approximation of particle velocities. Note that for one-dimensional open-boundary systems, the ideal probability current has to be zero at all positions.
@@ V Modified Temperature Profile and Probability Distribution
In Fig. 3, we find that, with the use of the effective temperature profile, the resulting NESS probability distribution of FPE analysis fits remarkably well with the one obtained by simulation. We have compared the simulation results with the FPE analysis using the effective temperature profile over various timesteps and temperature settings, while both results stay in pretty good agreement with each other.
The above comparison suggests that one can still derive the distribution of a one-dimensional stochastic simulation over discrete timesteps via an effective FPE analysis. However, this interpretation may appear to be paradoxical, since we noted earlier that any one-dimensional Fokker-Planck-type equation implies an underlying chain-connecting structure. And how could the detailed balance be violated at NESS if the underlying structure is chain-like with open boundaries? We speculate that this paradox may be resolved knowing that time irreversibility is implicitly incorporated through the effective temperature formalism. To be more specific, an effective temperature at a particular position contains the information of its last-step temperature, and this memory effect is in fact irreversible.
@@ Acknowledgements.
data availability
The data that support the findings of this article can be found in Ref. [28].